What is the final temperature when 150.0 g of water at 90.0oC is added to 100.0 g of water at 30.0 oC?


1
Expert's answer
2022-02-14T17:07:05-0500

Solution:

This problem can be summarized thusly:

-qlost by water1 = qgained by water2


q = m × C × ΔT

q = m × C × (Tf - Ti),

where:

q = amount of heat energy gained or lost by substance (J)

m = mass of sample (g)

C = specific heat capacity (J °C-1 g-1)

Tf = final temperature (°C)

Ti = initial temperature (°C)


The specific heat capacity of water is 4.184 J °C-1 g-1

 

Therefore:

qlost by water1 = (150.0) × (4.184) × (Tf - 90.0)

qlost by water1 = 627.6 × (Tf - 90.0)

 

qgained by water2 = (100.0) × (4.184) × (Tf - 30.0)

qgained by water2 = 418.4 × (Tf - 30.0)

 

-qlost by water1 = qgained by water2

-627.6 × (Tf - 90.0) = 418.4 × (Tf - 30.0)

-1.5 × (Tf - 90.0) = Tf - 30.0

-1.5 × Tf + 135 = Tf - 30.0

2.5 × Tf = 165

Tf = 66°C


Answer: The final temperature (Tf) is 66°C